# “It Explains What I am Currently Going Through Perfectly to a Tee”: Understanding User Perceptions on LLM-Enhanced Narrative Interventions

### v1

@@ 3. Study Design
Below, we describe the design of our study and the various steps we employed to execute it.
@@ 3.1. Procedure
Participants were recruited for a brief activity, lasting between 10 and 20 minutes, aimed at aiding the development of technology-based narrative interventions for managing negative emotions. During recruitment, it was communicated that their participation would contribute to research in this area. The survey activity, which included dynamic story selection and generation through the use of LLMs, was conducted on the Qualtrics platform.
@@
After providing affirmative informed consent, participants were asked to select a common challenging thought that they frequently encounter (we label this question as Q-Challenging-Thought). This selection was made from six options identified from previous research on mental health challenges in young adults (Meyerhoff et al., 2024). The options included feelings of being down on oneself, worrying about the future, feelings of struggling, feelings of underachievement compared to others, disappointments in reliance on others, and social anxiety. Participants then evaluated their belief in, and the emotional intensity of, the selected thought on a scale from 1 (’very low’) to 7 (’very high’), inspired by questions from a past study (Sharma et al., 2023b). They were also requested to provide a detailed description of their thought, circumstances, and emotions to the extent they felt comfortable (we label this question as Q-Challenging-Thought-Description).
@@
Following this initial sequence of questions, participants were randomly assigned to engage with either a story sourced from prior studies (Non-LLM condition) or an LLM-enhanced version of the story (LLM condition). In both groups, stories addressed the same challenging thought endorsed by the participant. Participants were blind to the condition they were assigned to; in both conditions, stories were described broadly as depicting the experience of someone else going through a similar situation. Each story concluded with a reflection question, prompting participants to reflect on their behavior and potential strategies to overcome their challenges. The number of words written by participants in this reflection was captured as an indicator of their engagement with the intervention (Wang et al., 2023). We detail the design of stories in both conditions in Section 3.3.
@@
Subsequently, participants rated four qualities of the stories — such as authenticity and the effectiveness of communicating key takeaways — on a scale from 1 (’very low’) to 7 (’very high’). They also provided qualitative feedback on their ratings and additional quantitative evaluations concerning related qualities like perceived relatability and engagement. Qualitative comments on both helpful and unhelpful aspects of the stories were also collected. These questions were also inspired from prior literature (Bhattacharjee et al., 2022, 2024a; Sharma et al., 2023b). Following best practices in survey design, we also incorporated attention-check questions in the middle of the survey to ensure participant attentiveness and the quality of the responses (Hauser and Schwarz, 2016).
@@
Table 1. Quantitative questions from the survey
Questions asked before and after reading the story
Question Label Actual Question Corresponding RQ Source
Belief in Negative Thought How strongly do you believe in your thought? RQ2 (Sharma et al., 2023b)
Emotion Intensity How strong is the emotion associated with this thought? RQ2 (Sharma et al., 2023b)
Questions asked after reading the story
Question Label Actual Question Corresponding RQ Source
Perceived Authenticity How would you rate the authenticity of the story in portraying characters, settings, and emotions? RQ1 (Bhattacharjee et al., 2022)
Perceived Ability to Communicate Key Takeaways How effectively do you think the story communicated its key messages or lessons? RQ1 (Bhattacharjee et al., 2022)
Perceived Ability to Promote Reflection How much did the story encourage you to reflect on your own thoughts, experiences, or beliefs? RQ1 (Bhattacharjee et al., 2022)
Perceived Balance of Positivity with Realistic Struggle How well did the story balance positive outcomes with realistic portrayals of challenges? RQ1 (Bhattacharjee et al., 2022)
Perceived Relatability Agreement with the statement: ‘I could relate to the experiences shared in the story.’ RQ2 (Meyerhoff et al., 2024)
Perceived Capacity of Inspiring Solutions Agreement with the statement: ‘The story helped me think of potential ways I can manage negative thoughts.’ RQ2 (Bhattacharjee et al., 2024a)
Likelihood of Future Engagement Agreement with the statement: ‘I would love to read similar stories in future.’ RQ2 (Bhattacharjee et al., 2024a)
For our quantitative comparison, we hypothesized that participants would perceive the LLM-enhanced stories more favorably across all qualities except for authenticity, compared to those in the Non-LLM condition. This hypothesis was based on the LLMs’ potential to improve DMH content across various settings. However, we expected lower scores for authenticity in the LLM condition because these stories, while based on human-written content, were adapted by LLMs, unlike the stories in the Non-LLM condition.
@@ 3.2. Participants
A total of 180 crowdworkers and 179 students initially participated in our study. However, data from 6 crowdworkers and 7 students were discarded due to quality issues identified through instructional manipulation checks of attention-check questions  (Oppenheimer et al., 2009), resulting in a final sample of 174 crowdworkers and 172 students.
@@
Table 2. Demographics of crowdworkers and students
Crowdworkers (N = 174) Students (N = 172)
Age, mean ± std err 22.5 ± 0.2 20.9 ± 0.1
Gender, N (%)
Women 101 (58.0%) 25 (14.5%)
Men 63 (36.2%) 142 (82.6%)
Non-binary/third gender 7 (4.0%) 0 (0%)
Self-described 1 (0.6%) 2 (1.2%)
Not reported 2 (1.1%) 3 (1.7%)
Race, N (%)
White 68 (39.1%) 19 (11.0%)
More than one race 13 (7.5%) 6 (3.5%)
Black or African American 13 (7.5%) 6 (3.5%)
Asian 65 (37.4%) 126 (73.3%)
American Indian or Alaskan Native 2 (1.1%) 0 (0%)
Not reported 13 (7.5%) 15 (8.7%)
3.3. Design of the Stories
In this section, we explain the design of stories in the Non-LLM and the LLM conditions. All of the stories in the Non-LLM condition, accompanied by an example LLM-enhanced version, are provided in Appendix A.
3.3.1. Design of Stories in the Non-LLM Condition
In the Non-LLM condition, our stories are sourced from Meyerhoff et al. (2024), who themselves based their stories on the experiences reported by young adult participants according to Kornfield et al. (2022a). To generate stories, Crowdworkers were given a list of common challenges to write about, or they could supply their own. They were then asked how they took a step that helped them better manage the challenge or feel better about it, even if it was a small or incremental step. They were provided with examples of evidence-based psychological strategies they may have used in addressing their challenge, or they were allowed to describe their own approach. Crowdworkers were asked to write true stories in the first person and to include as much detail as they could about their challenge and how they had addressed it, all while omitting identifying information. After clinical psychology and human-computer interaction researchers edited and curated the stories for clarity and to align with the intended psychological strategies, Meyerhoff et al. (2024) eventually deployed these stories as a part of a text messaging program designed to help young adults manage mental health and wellbeing.
@@ 3.3.1. Design of Stories in the Non-LLM Condition
3.3.2. Design of Stories in the LLM Condition
Our implementation of LLM-enhanced narratives was informed by recent research across various contexts, which demonstrated the benefits of adapting content to align with user contexts. (Bhattacharjee et al., 2024b; Stade et al., 2024). For our project, we employed a prompt engineering method. The prompt design was an iterative effort involving faculty and graduate students with expertise in human-computer interaction, LLMs, psychology, cognitive science, and DMH. The prompt design underwent multiple rounds of revision, applying the prompts to a variety of stories from past interventions and refining them iteratively.
Following recommended practices of prompt engineering (Stade et al., 2024; Ekin, 2023), we developed a structured prompt using OpenAI’s GPT-4, incorporating example stories from the Non-LLM condition to ensure adherence to the structured format discussed in Section 3.3.1. The prompt was designed to weave in three responses from participants:
  • Selected option from Q-Challenging-Thought
  • Response to the open-ended question Q-Challenging-Thought-Description
  • Corresponding story to the selected option from Q-Selected-Thought
We also included instructions to avoid the introduction of new suggestions or strategies not present in the examples. The stories were designed to maintain a similar length to those in the Non-LLM condition ($\leq$250 words) and concluded with a reflective question. The complete prompt is as follows:
@@ 3.3.2. Design of Stories in the LLM Condition
[backgroundcolor=gray!20]Given the user’s thought, [[selected choice from Q-Challenging-Thought]], and their provided description, [[response to the open-ended question Q-Challenging-Thought-Description]], your task is to share a narrative. This narrative should recount the journey of an individual who encountered challenges similar to yours and the steps they took towards making progress, incorporating relevant examples and strategies for improvement. Here are some guidelines to follow:
@@
Please generate a first-person narrative that incorporates these guidelines and examples, making it as engaging and helpful as possible for the user to gain new perspectives and coping mechanisms. The text should start with “Here’s someone’s experience that might resonate with you: [[newline]]” and should not contain more than 250 words. As shown in the example, conclude with a clear separator, “******,” followed by a prompt for a reflective question that starts with “Please answer this reflective question:”
@@ 3.4. Data Analysis
After the data were anonymized and cleaned, we utilized the thematic analysis approach outlined by Clarke and Braun (2017) to examine the qualitative data. First, two coders from the research team reviewed the data to familiarize themselves with the data. Using an open-coding method (Khandkar, 2009), each coder independently developed an initial codebook for the data of 25 participants. We then implemented a consensus coding method similar to that outlined by Clarke and Braun (2017). During this phase, the coders held several joint sessions to review and integrate their separate codebooks. This consensus coding process, through repeated discussions and refinement, was designed to improve reliability and minimize personal biases. The coders determined common codes, refined the definitions, and removed codes that did not directly relate to the research questions. The updated codebook was then applied iteratively to the data from an additional 25 participants, making adjustments as new findings emerged. After finalizing the codebook, each coder independently applied the established codes to half of the remaining data.
Our analysis generated 138 codes, capturing themes such as “new perspective,” “matching personal struggles,” “robotic,” and “clear takeaways.” Coders used axial coding (Williams and Moser, 2019) in subsequent meetings to organize these codes into major themes. Detailed findings structured by these themes are presented in Section 4.
@@ 3.5. Ethical Considerations
Moreover, while we did not ask participants about their experiences with suicide-related thoughts and behaviors, we recognized that suicidal thoughts, behaviors, or other risk-related issues may arise due to the open-ended nature of the questions and assessments as well as the probabalistic nature of LLM responses. We provided all participants with contact information for crisis services, such as suicide prevention helplines, to ensure they had access to support, if needed. Additionally, the research team reviewed all survey responses daily during the study period to monitor for any indications of suicide-related risk, such as suicidal ideation or self-harm. Team members were trained to respond appropriately, and assess participants using the Columbia-Suicide Risk Assessment protocol (Posner et al., 2008) if any potential suicide or self-harm-related responses were identified. However, no such risks emerged during the course of the study, and we did not need to conduct any follow-up assessments.
@@ 4. Results
We structure our findings around the research questions we originally set. First, we examine how LLM-enhanced stories were perceived in terms of exhibiting desirable narrative qualities. Subsequently, we explore their effectiveness as interventions, specifically looking at their impact on managing negative thoughts. We primarily explore user perceptions of stories in the LLM condition, occasionally drawing comparisons to the Non-LLM condition to highlight differences and similarities.
4.1. LLM-Enhanced Stories in Exhibiting Key Narrative Qualities
Table 3 presents the summary statistics of the ratings given by participants regarding the narrative qualities of the stories. We observed that the stories from the LLM condition were perceived to be superior in all the narrative qualities other than authenticity. The differences between groups in authenticity ratings were not significant. The remaining differences were statistically significant among the crowdworkers, whereas among the students, only the difference in the perceived ability to promote reflection was significant. Below, we support these trends using insights drawn from participants’ comments.

### v4

@@ 3. Study Design
We describe the design of our study and the various steps we employed to execute it below.
@@ 3.1. Procedure
Participants were recruited for a brief activity, lasting between 10 and 20 minutes, aimed at aiding the development of technology-based narrative interventions for managing negative thoughts and emotions. The study targeted members of the general population in order to explore diverse experiences with common negative thoughts and emotions. During recruitment, it was communicated that their participation would contribute to research in this area. The study activity, which included dynamic story selection and generation through the use of LLMs, was conducted on the Qualtrics platform.
@@
After providing affirmative informed consent, participants were asked to select a common challenging thought that they frequently encounter (we label this question as Q-Challenging-Thought). This selection was made from six options identified from previous research on mental health challenges in young adults (Meyerhoff et al., 2024). The options included feelings of being down on oneself, worrying about the future, feelings of struggling, feelings of underachievement compared to others, disappointments in reliance on others, and social anxiety. Participants then evaluated their belief in, and the emotional intensity of, the selected thought on a scale from 1 (‘very low’) to 7 (‘very high’), inspired by questions from a past study (Sharma et al., 2023b). They were also requested to provide a detailed description of their thought, circumstances, and emotions to the extent they felt comfortable (we label this question as Q-Challenging-Thought-Description).
@@
Following this initial sequence of questions, participants were randomly assigned to engage with either a story sourced from prior studies (Non-LLM condition) or a personalized LLM-enhanced version of the story (LLM condition). In both groups, stories addressed the same challenging thought endorsed by the participant. Participants were blinded to their assigned condition in that the story was broadly described as illustrating how someone might navigate a situation similar to the one the participant had mentioned. Each story concluded with a reflection question, prompting participants to reflect on their behavior and potential strategies to overcome their challenges. The number of words written by participants in this reflection was captured as an indicator of their engagement with the intervention (Wang et al., 2023). We detail the design of stories in both conditions in Section 3.3.
@@
Subsequently, participants rated four qualities of the stories — such as authenticity and the effectiveness of communicating key takeaways — on a scale from 1 (‘very low’) to 7 (‘very high’). They also provided qualitative feedback on their ratings and additional quantitative evaluations concerning related qualities like perceived relatability and engagement. Qualitative comments on both helpful and unhelpful aspects of the stories were also collected. These questions were also inspired from prior literature (Bhattacharjee et al., 2022, 2024a; Sharma et al., 2023b). Following best practices in survey design, we also incorporated attention-check questions in the middle of the activity to ensure participant attentiveness and the quality of the responses (Hauser and Schwarz, 2016).
Table 1. Quantitative questions from the survey
[…]
Quantitative survey questions used to evaluate the stories. Includes pre- and post-story measures such as belief in negative thought and emotion intensity, as well as post-story assessments of narrative qualities such as authenticity, takeaways, reflection, relatability, and future engagement. Each item is linked to a corresponding research question and literature source. Questions asked before and after reading the story Question Label Actual Question Corresponding RQ Source Belief in Negative Thought How strongly do you believe in your thought? RQ2 (Sharma et al., 2023b) Emotion Intensity How strong is the emotion associated with this thought? RQ2 (Sharma et al., 2023b) Questions asked after reading the story Question Label Actual Question Corresponding RQ Source Perceived Authenticity How would you rate the authenticity of the story in portraying characters, settings, and emotions? RQ1 (Bhattacharjee et al., 2022) Perceived Ability to Communicate Key Takeaways How effectively do you think the story communicated its key messages or lessons? RQ1 (Bhattacharjee et al., 2022) Perceived Ability to Promote Reflection How much did the story encourage you to reflect on your own thoughts, experiences, or beliefs? RQ1 (Bhattacharjee et al., 2022) Perceived Balance of Positivity with Realistic Struggle How well did the story balance positive outcomes with realistic portrayals of challenges? RQ1 (Bhattacharjee et al., 2022) Perceived Relatability Agreement with the statement: ‘I could relate to the experiences shared in the story.’ RQ2 (Meyerhoff et al., 2024) Perceived Capacity of Inspiring Solutions Agreement with the statement: ‘The story helped me think of potential ways I can manage negative thoughts.’ RQ2 (Bhattacharjee et al., 2024a) Likelihood of Future Engagement Agreement with the statement: ‘I would love to read similar stories in future.’ RQ2 (Bhattacharjee et al., 2024a)
@@
For our quantitative comparison, we hypothesized that participants would perceive the personalized LLM-enhanced stories more favorably across all qualities except for authenticity, compared to those in the Non-LLM condition. This hypothesis was based on the LLMs’ potential to improve DMH content across various settings. However, we expected lower scores for authenticity in the LLM condition because these stories, while based on human-written content, were adapted by LLMs, unlike the stories in the Non-LLM condition.
Although this was not a clinical intervention, the study was designed and carried out in close collaboration with licensed clinical psychologists who were part of the research team throughout the project. They contributed to all major aspects of the intervention, including the creation of stories in the Non-LLM condition (Section 3.3.1), the design of prompts for the LLM condition (Section 3.3.2), and the implementation of ethical considerations (Section 3.5).
@@ 3.2. Participants
A total of 180 crowdworkers and 179 students initially participated in our study. However, data from 6 crowdworkers and 7 students were discarded due to quality issues identified through instructional manipulation checks (Oppenheimer et al., 2009), resulting in a final sample of 174 crowdworkers and 172 students.
@@
Table 2. Demographics of the crowdworkers and students
[…]
Demographic characteristics of participants, including mean age, gender distribution, and racial identity for both crowdworkers and students. Highlights demographic differences between the two populations. Crowdworkers (N = 174) Students (N = 172) Age, mean ± std err 22.5 ± 0.2 20.9 ± 0.1 Gender, N (%) Women 101 (58.0%) 25 (14.5%) Men 63 (36.2%) 142 (82.6%) Non-binary/third gender 7 (4.0%) 0 (0%) Self-described 1 (0.6%) 2 (1.2%) Not reported 2 (1.1%) 3 (1.7%) Race, N (%) White 68 (39.1%) 19 (11.0%) More than one race 13 (7.5%) 6 (3.5%) Black or African American 13 (7.5%) 6 (3.5%) Asian 65 (37.4%) 126 (73.3%) American Indian or Alaskan Native 2 (1.1%) 0 (0%) Not reported 13 (7.5%) 15 (8.7%)
3.3. Design of the Intervention in the Study
In this section, we describe the design of our intervention in both the Non-LLM and LLM conditions. In both cases, the intervention was implemented as an SSI—a structured, standalone session designed to provide focused support with immediate effects on specific outcomes. SSIs can be highly effective in delivering rapid relief and promoting reflection or reappraisal within a short timeframe (typically 10-20 minutes), making them accessible and adaptable to diverse populations (Schleider and Weisz, 2017; Schleider et al., 2020). They can also serve as an entry point to broader mental health care by helping individuals develop insights or skills that may encourage further engagement with mental health resources (Bhattacharjee et al., 2024a). Research has demonstrated their success in managing negative thoughts, reducing stress, and alleviating symptoms of anxiety and depression in both clinical and non-clinical settings (Sharma et al., 2023b; Schleider and Weisz, 2017; Schleider et al., 2020; Bhattacharjee et al., 2024a).
Building on prior SSIs for managing negative thoughts (Sharma et al., 2023b; Bhattacharjee et al., 2024a) and leveraging the potential of narratives to assist individuals in addressing negative thoughts and situational challenges (Bhattacharjee et al., 2022; Meyerhoff et al., 2024), we developed an SSI centered on using stories to support participants in managing common negative thoughts. Stories in the Non-LLM condition, along with an example LLM-enhanced version for each, are included in Appendix A. It is important to note that, even when using the same negative thoughts and descriptions as inputs, the LLM may generate outputs that differ from the provided examples. We provide a detailed explanation of the design of stories below.
3.3.1. Design of the Stories in the Non-LLM Condition
In the Non-LLM condition, our stories were sourced from Meyerhoff et al. (2024), who themselves based their stories on the experiences reported by young adult participants according to Kornfield et al. (2022a). To generate stories, a separate group of crowdworkers distinct from those who participated in the main study were given a list of common challenges to write about, or they were allowed to supply their own. They were then asked how they took a step that helped them better manage the challenge or feel better about it, even if it was a small or incremental step. They were provided with examples of evidence-based psychological strategies they may have used in addressing their challenge, or they were allowed to describe their own approach. Crowdworkers were asked to write true stories in the first person and to include as much detail as they could about their challenge and how they had addressed it, all while omitting identifying information. After clinical psychology and human-computer interaction researchers edited and curated the stories for clarity and to align with the intended psychological strategies, Meyerhoff et al. (2024) eventually deployed these stories as a part of a text messaging program designed to help young adults manage mental health and wellbeing.
@@ 3.3.1. Design of the Stories in the Non-LLM Condition
3.3.2. Design of the Stories in the LLM Condition
Our implementation of personalized LLM-enhanced narratives was informed by recent research across various contexts, which demonstrated the benefits of adapting content to align with user contexts (Bhattacharjee et al., 2024c; Stade et al., 2024). For our project, we employed a prompt engineering method. The prompt design was an iterative effort involving faculty and graduate students with expertise in human-computer interaction, LLMs, psychology, cognitive science, and DMH. The prompt design underwent multiple rounds of revision, applying the prompts to a variety of stories from past interventions and refining them iteratively. Following recommended practices of prompt engineering (Stade et al., 2024; Ekin, 2023), we developed a structured prompt using OpenAI’s GPT-4 that aimed to achieve the following objectives:
  • Dynamic Personalization: We aimed to generate stories during the interaction that reflect participants’ unique situations and challenges. This was achieved by incorporating two responses from participants: the selected option from Q-Challenging-Thought and the response to the open-ended question Q-Challenging-Thought-Description.
  • Adherence to Guidelines: We aimed to ensure that the generated narratives align with the intended qualities in prior DMH interventions. This also involved integrating prior relevant examples by including the corresponding story to the selected option from Q-Selected-Thought. We included instructions to avoid the introduction of new suggestions or strategies not present in the examples. The stories were designed to maintain a similar length to those in the Non-LLM condition ($\leq$250 words) and concluded with a reflective question.
The complete prompt is as follows:
@@ 3.3.2. Design of the Stories in the LLM Condition
[backgroundcolor=gray!20]Given the user’s thought, [[selected choice from Q-Challenging-Thought]], and their provided description, [[response to the open-ended question Q-Challenging-Thought-Description]], your task is to share a narrative. This narrative should recount the journey of an individual who encountered challenges similar to user’s and the steps they took towards making progress, incorporating relevant examples and strategies for improvement. Here are some guidelines to follow:
@@
Please generate a first-person narrative that incorporates these guidelines and examples, making it as engaging and helpful as possible for the user to gain new perspectives and coping mechanisms. The text should start with “Here’s someone’s experience that might resonate with you: [[newline]]” and should not contain more than 250 words. As shown in the example, conclude with a clear separator, “******,” followed by a prompt for a reflective question that starts with “Please answer this reflective question:”.
@@ 3.4. Data Analysis
After the data were anonymized and cleaned, we utilized the thematic analysis approach outlined by Clarke and Braun (2017) to examine the qualitative data. First, two coders from the research team reviewed the data to familiarize themselves with the data. Using an open-coding method (Khandkar, 2009), each coder independently developed an initial codebook for the data of 25 participants. This process reflected an iterative approach that integrated both deductive and inductive techniques. The analysis of RQ1 was guided by prior work on narrative qualities (Bhattacharjee et al., 2022) and followed a deductive coding approach, whereas RQ2 was approached inductively, allowing themes to emerge directly from the data. We then implemented a consensus coding method similar to that outlined by Clarke and Braun (2017). During this phase, the coders held several joint sessions to review and integrate their separate codebooks. This consensus coding process, through repeated discussions and refinement, was designed to improve reliability and minimize personal biases. The coders determined common codes, refined the definitions, and removed codes that did not directly relate to the research questions. The updated codebook was then applied iteratively to the data from an additional 25 participants, making adjustments as new findings emerged. After finalizing the codebook, each coder independently applied the established codes to half of the remaining data.
Our analysis generated 138 codes such as “new perspective”, “matching personal struggles”, “giving hope”, “detailed character portrayal”, “robotic”, “informative”, and “clear takeaways”. Coders used axial coding (Williams and Moser, 2019) in subsequent meetings to organize these codes into eight major themes (four per RQ). Detailed findings structured by these themes are presented in Section 4.
@@ 3.5. Ethical Considerations
Moreover, while we did not specifically ask participants about their experiences with suicide-related thoughts and behaviors, we recognized that suicidal thoughts, behaviors, or other risk-related issues may arise due to the open-ended nature of the questions and assessments as well as the probabalistic nature of LLM responses. We provided all participants with contact information for crisis services, such as suicide prevention helplines, to ensure they had access to support, if needed. Additionally, the research team reviewed all stories and participant responses daily during the study period to monitor for any indications of suicide-related risk (e.g., suicidal ideation, self-harm). Team members were trained to assess participants using the Columbia-Suicide Risk Assessment protocol (Posner et al., 2008) if any potential suicide or self-harm-related responses were identified. However, no such risks emerged during the course of the study, and we did not need to conduct any follow-up assessments.
@@ 4. Results
We structure our findings around the research questions we originally set. First, we examine how personalized LLM-enhanced stories were perceived in terms of exhibiting desirable narrative qualities. Subsequently, we explore their effectiveness as interventions, specifically looking at their impact on managing negative thoughts. We primarily explore user perceptions of stories in the LLM condition, occasionally drawing comparisons to the Non-LLM condition to highlight differences and similarities.
4.1. Personalized LLM-Enhanced Stories in Exhibiting Key Narrative Qualities
Table 3 presents the summary statistics of the ratings given by participants regarding the narrative qualities of the stories. Additional details about ratings across different challenging thoughts are provided in Appendix B. We observed that the stories from the LLM condition were perceived to be superior in all the narrative qualities other than authenticity. The differences between groups in authenticity ratings were not significant. The remaining differences were statistically significant among the crowdworkers, whereas among the students, only the difference in the perceived ability to promote reflection was significant. Below, we support these observations using insights drawn from participants’ comments.